# A Generalized Sylvester-Fermat-Torricelli problem with application in disaster relief operations by UAVs

### v1

@@ 6.3 Application of the Generalized SFT Problem
$\min_{x\in\Omega_{0}}S(x),$ (6.14)
where
$S(x):=\max\{\rho_{F_{i}}^{\Omega_{i}}(x)+\rho_{\bar{F}_{i}}^{\Omega_{i}}(x)\ |\ i=1,2,3\},$ (6.15)
such that, $\bar{F}_{i}=-F_{i}$.
To compute $\partial S(x)$, we apply Theorems 2.3 and 2.4, yielding:
@@
Figure 4: Case 1 in Table 3: (a) Approximated optimal location of the truck when the wind effect is neglected. (b) Dynamic sets with $\|\mathcal{S}\|=0$. (c) Approximated optimal location of the truck under windy conditions. (d) Dynamic sets of the UAVs under windy conditions with $\mathcal{S}=(-0.6,0.6)$. The black arrow indicates the wind direction
@@
where $\partial\rho_{F_{i}}^{\Omega_{i}}(x)$ and $\partial\rho_{\bar{F}_{i}}^{\Omega_{i}}(x)$ is obtained by procedure outlined in Section 6.1. The subgradient algorithm (Algorithm 1) can be utilized to compute an optimal solution to problem (6.14). In continuation, we consider a more challenging case of this scenario, which has a real-world application. In this case, UAVs are deployed from a point $x$ (the location of the mobile station), first travel to the nearest point in $\Omega_{i}$, and then return from the farthest point in $\Omega_{i}$ back to $x$. Such cases commonly arise when UAVs are used to cover regions $\Omega_{i}$ for tasks such as data collection or aerial photography of affected areas, employing coverage path planning algorithms (see, for example, [13, 5, 4, 16]). Then, in the worst-case situation, UAVs complete their coverage at the farthest point and then must return to the mobile station for recharging and data transfer. This case can be formulated through the extended version of the generalized SFT problem (5.1), as follows:
$\min_{x\in\Omega_{0}}\hat{S}(x),$ (6.17)
where
$\hat{S}(x):=\max\{\rho_{\bar{F}_{i}}^{\Omega_{i}}(x)+\bar{\rho}_{F_{i}}^{\Omega_{i}}(x)\ |\ i=1,2,3\}$ (6.18)
such that $\bar{F}_{i}=-F_{i}$.
To obtain $\partial S(x)$, we apply Theorems 2.3 and 2.4, yielding:
@@
$\partial\hat{S}(x)=\operatorname{co}\left\{\bigcup_{i\in I(x)}\left(\partial\rho_{\bar{F}_{i}}^{\Omega_{i}}(x)+\partial\bar{\rho}_{F_{i}}^{\Omega_{i}}(x)\right)\,\middle|\,i=1,2,3\right\}.$ (6.19)
@@
We use Algorithm 1 to find an approximate minimizer of this problem. At each step of the algorithm, a subgradient from $\partial\hat{S}(x)$ is required. To obtain this, we compute $\partial\rho_{\bar{F}_{i}}^{\Omega_{i}}(x)$ by following the procedure described in Section 6.1. Moreover, for simplicity, instead of fully calculating $\partial\bar{\rho}_{F_{i}}^{\Omega_{i}}(x)$, we employ its estimation provided in Theorem 3.9. In particular, we have
@@
$\nabla\rho_{F_{i}}(x-\omega^{\prime}_{i})\subseteq\partial\bar{\rho}_{F_{i}}^{\Omega_{i}}(x),$
where $\omega_{i}^{\prime}\in\bar{\Pi}_{F_{i}}(x;\Omega_{i})$, as defined in Theorem 3.9. Finally, by (6.19), obtaining a subgradient of $\partial\hat{S}(x)$ becomes straightforward.
For the problem in (6.17), associated with Case 1 in Table 3, the first step is to identify the approximate optimal location of the truck under the condition $\|\mathcal{S}\|=0$, with the corresponding dynamic sets shown in Fig. 4(b). The obtained solution is $(171,134)$, as shown in Fig. 4(a), with a corresponding transition time of $905\,\text{s}$ under real conditions with wind. When incorporating the wind vector $\mathcal{S}=(-0.6,0.6)$, the dynamic sets change accordingly, as illustrated in Fig. 4(d). In this case, the approximate optimal location of the truck shifts to $(189,56)$, as shown in Fig. 4(c), with a reduced transition time of $401\,\text{s}$. This represents a $55.6\%$ improvement, demonstrating the effectiveness of the generalized SFT problem (extended variant) in addressing weather uncertainty. These results highlight that neglecting wind effects can result in a time loss of nearly $500\,\text{s}$—an inefficiency that may be critical in time-sensitive rescue operations.

### v2

@@ 6.3 Applications of the Generalized SFT Problem
$\min_{x\in\Omega_{0}}S(x),\;\mbox{ where }\;S(x):=\max\big\{\rho_{F_{i}}^{\Omega_{i}}(x)+\rho_{\overline{F}_{i}}^{\Omega_{i}}(x)\ \big|\ i=1,2,3\big\}.$ (6.14)
Assuming that $\overline{F}_{i}=-F_{i}$, compute $\partial S(x)$ by applying Theorems 2.3 and 2.4. This gives us
@@
Figure 5: Case 1 in Table 3: (a) Approximated optimal location of the truck when the wind effect is neglected. (b) Dynamic sets with $\|\mathcal{S}\|=0$. (c) Approximated optimal location of the truck under windy conditions. (d) Dynamic sets of the UAVs under windy conditions with $\mathcal{S}=(-0.6,0.6)$. The black arrow indicates the wind direction.
@@
where $\partial\rho_{F_{i}}^{\Omega_{i}}(x)$ and $\partial\rho_{\overline{F}_{i}}^{\Omega_{i}}(x)$ are calculated similarly to Section 6.1. The subgradient algorithm (Algorithm 1) can be utilized to compute an optimal solution to problem (6.14). In continuation, we consider a more challenging case of this scenario, which has a real-world application. In this case, UAVs are deployed from a point $x$ (the location of the mobile station), first travel to the nearest point in $\Omega_{i}$, and then return from the farthest point in $\Omega_{i}$ back to $x$. Cases of this type commonly arise when UAVs are used to cover regions $\Omega_{i}$ for tasks such as data collection or aerial photography of affected areas, employing coverage path planning algorithms (see, for example, [kazemdehbashi2025algorithm, Coombes2018, Choset2000, Latombe1991]). Then, in the worst-case situation, UAVs complete their coverage at the farthest point and then must return to the mobile station for recharging and data transfer. This can be formulated via the extended version of the generalized SFT problem (5.2) as follows:
$\min_{x\in\Omega_{0}}\widehat{S}(x),\;\mbox{ where }\;\widehat{S}(x):=\max\big\{\rho_{\overline{F}_{i}}^{\Omega_{i}}(x)+\overline{\rho}_{F_{i}}^{\Omega_{i}}(x)\ \big|\ i=1,2,3\big\}.$ (6.16)
Assuming again that $\overline{F}_{i}=-F_{i}$ gives us by Theorems 2.3 and 2.4 that
@@
$\partial\widehat{S}(x)=\operatorname{co}\left\{\bigcup_{i\in I(x)}\left(\partial\rho_{\overline{F}_{i}}^{\Omega_{i}}(x)+\partial\overline{\rho}_{F_{i}}^{\Omega_{i}}(x)\right)\,\middle|\,i=1,2,3\right\}.$ (6.17)
@@
Now we use Algorithm 1 to find an approximate minimizer of this problem. At each step of the algorithm, a subgradient from $\partial\widehat{S}(x)$ is required. To this end, we compute $\partial\rho_{\overline{F}_{i}}^{\Omega_{i}}(x)$ similarly to Section 6.1. To simplify the procedure, the lower estimate of $\partial\overline{\rho}_{F_{i}}^{\Omega_{i}}(x)$ from Theorem 3.9 can be used instead of the full calculation of the subdifferential. In this way, we get
@@
$\nabla\rho_{F_{i}}(x-\omega^{\prime}_{i})\subset\partial\overline{\rho}_{F_{i}}^{\Omega_{i}}(x)\;\mbox{ with }\;\omega_{i}^{\prime}\in\overline{\Pi}_{F_{i}}(x;\Omega_{i})$
Then determining a subgradient of $\partial\widehat{S}(x)$ becomes straightforward from (6.17).
Figure 6: (a) Effect of wind direction on transition time at approximately similar wind speeds. The gray line shows transition time without considering wind, while the yellow line reflects results from the generalized SFT model accounting for wind. Results correspond to Cases 14–21 in Table 3. (b) Impact of wind vector norm (speed) on transition time. Results correspond to Cases 1–7 in Table 2.
a) Refer to caption   b) Refer to caption
If in the case modeled by problem (6.16), UAVs are deployed at the earliest time and the objective is to determine an optimal location of the mobile station for collecting the data as described in Section 6.2, the extended generalized Sylvester problem (5.6) can be used to optimize the location of the mobile station. Furthermore, if the rescue team employs a single UAV instead of three (as in the scenario presented in Section 6.1), the UAV travels from the mobile station to the closest point of $\Omega_{i}$ and returns from the farthest point while repeating this pattern for the subsequent reference sets. The extended generalized Fermat-Torricelli problem (5.4) can then be used to determine the optimal location of the mobile station.
Table 3 summarizes 24 cases with different parameter settings to evaluate the efficiency of the generalized SFT problem (extended variant) in accounting for wind effects. For simplicity, all cases are conducted on a small scale, with the slowest UAV assigned a nominal speed of 11(m/s); these values can be appropriately scaled to represent real-world scenarios. The results show that the generalized SFT problem (extended variant) reduces the transition time by approximately $66.8\%$. Furthermore, Cases 14–21 share identical settings except for the direction of the wind vector. Fig. 6(a) illustrates these results that confirm the ability of the generalized SFT formulation to address wind effects across varying directions.